# Axion-like Dark Matter Search with Space-based Gravitational Wave Detectors

### v1

@@ III Proposed Modification
$\Delta\phi=\Delta\phi_{\mathrm{Tx}}+\Delta\phi_{\mathrm{Rx}}+\Delta\phi_{\mathrm{PS}}$ (19)
represents the phase shift relative to the ISI. Here, $\Delta\phi_{\mathrm{Tx}}$ and $\Delta\phi_{\mathrm{Rx}}$ are the phase differences arising from the path length differences of the transmitted light and received light, respectively. $\Delta\phi_{\mathrm{PS}}$ is the phase shift introduced by the phase shifter, which can be preset on Earth. The term $I_{1}$ is the power of the reference beam in the ALDM interferometer, and $p_{\mathrm{ALDM}}^{2}=p_{C}^{2}+p_{D}^{2}$ and $u_{\mathrm{ALDM}}^{2}=u_{C}^{2}+u_{D}^{2}$ denote the ratios of the power of the received light and the backscattered light to the reference beam, respectively. Likewise, the power of the received light in the ALDM interferometer is modulated by the birefringence effect
$p_{\mathrm{ALDM}}^{\prime}=\sin\Delta\phi_{a}p_{\mathrm{ALDM}}.$ (20)
For simplicity, we assume $I_{1}=I_{0}$, which implies $p_{\mathrm{ALDM}}=p_{\mathrm{ISI}}$. In addition, the heterodyne signals are considered to be balanced Carleton and Maloney (1968); Fleddermann et al. (2018) with $p_{k}=p/\sqrt{2},\,k\in\{A,B,C,D\}$. Under these conditions, the combined AC components of the signal differences from both interferometers can be expressed as
$AC\{(S_{D}-S_{C})+(S_{B}-S_{A})\}$ (21)
$=$ $\sqrt{2}pI_{0}\Big{[}\cos\Delta\phi_{a}\sin(\Omega t+\phi)+\sin\Delta\phi_{a}\sin(\Omega t+\phi+\Delta\phi)$
$+$ $(u_{B}-u_{A})\cos(\Omega t+\phi-\psi)+(u_{D}-u_{C})\cos(\Omega t+\phi+\Delta\phi-\tilde{\psi})\Big{]}$
$=$ $\sqrt{2}pI_{0}\Big{[}\sin(\Omega t+\phi+\Delta\phi_{a})+\sin\Delta\phi_{a}[\sin(\Omega t+\phi+\Delta\phi)-\cos(\Omega t+\phi)]$
$+$ $(u_{B}-u_{A})\cos(\Omega t+\phi-\psi)+(u_{D}-u_{C})\cos(\Omega t+\phi+\Delta\phi-\tilde{\psi})\Big{]}.$
@@
$AC\{(S_{D}-S_{C})+(S_{B}-S_{A})\}$ (22)
$=$ $\sqrt{2}pI_{0}\Big{[}\sin(\Omega t+\phi+\Delta\phi_{a})$
$+$ $(u_{B}-u_{A})\cos(\Omega t+\phi-\psi)-(u_{D}-u_{C})\sin(\Omega t+\phi-\tilde{\psi})\Big{]}.$
The choice of $\Delta\phi=\pi/2$ is a crucial, as it maximizes the sensitivity to the ALDM signal by ensuring that the auxiliary interferometer operates at the optimal phase difference relative to the main interferometer. This configuration converts the amplitude modulation into a phase shift that can be directly measured by standard phase detection techniques.
@@
$\phi_{\mathrm{ALDM}}\approx\phi+\Delta\phi_{a}+\pi/2+\phi_{b},$ (23)
@@
$\phi_{b}\approx\left[(u_{A}-u_{B})\cos\psi-(u_{C}-u_{D})\sin\tilde{\psi}\right]\left[1-(u_{A}-u_{B})\sin\psi-(u_{C}-u_{D})\cos\tilde{\psi}\right].$ (24)
The phase $\phi_{\mathrm{ALDM}}$ is the key measurand of the ALDM interferometer. By subtracting $\phi_{\mathrm{ISI}}$ from $\phi_{\mathrm{ALDM}}$, we can recover the change in the polarization angle induced by the ALDM for a single-link light, as expressed in Eq. (8).
@@
IV Sensitivity to the ALDM-photon coupling
@@
IV.1 Single-arm signal
By differentiating the phase difference induced by the ALDM-induced birefringence, we define the single-arm response as  [^1] The single-arm signal seems a little different from the equation in Gué et al. (2025), this originates from the typo of the definition of the reduced Planck energy in their paper, which is supposed to be $E_{P}=\sqrt{\hbar c/\kappa}\simeq 2.435\times 10^{18}\,\mathrm{GeV}$. With the correct definition, the factor $\sqrt{8\pi G}$ would be canceled out.
$\eta_{rs}$ $\equiv\frac{1}{\omega_{r}}\frac{d\left(\phi_{\mathrm{ALDM}}-\phi_{\mathrm{ISI}}\right)}{dt}\approx\frac{1}{\omega_{r}}\frac{d\Delta\phi_{a}}{dt}$ (25)
@@ IV.1 Single-arm signal
$\bar{\eta}_{rs}\equiv\frac{1}{\omega_{r}}\frac{d\left(\phi_{\mathrm{ISI}}-\phi_{\mathrm{ALDM}}\right)}{dt}\approx-\frac{1}{\omega_{r}}\frac{d\Delta\phi_{a}}{dt},$ (26)
where $\omega_{r}$ denotes the laser frequency, and the subscripts $r$ and $s$ represent the indices of the S/C that received and sent the light beam, respectively. The single-arm responses $\eta_{rs}$ and $\bar{\eta}_{rs}$ provide a direct measure of the time derivative of the ALDM-induced phase shift. In practice, SGWDs employ TDI to suppress noises, such as laser frequency noise and OB motion noise, by combining the signals from multiple interferometric arms with appropriate time delays Armstrong et al. (1999); Tinto and Armstrong (1999); Dhurandhar et al. (2002); Tinto and Dhurandhar (2021). This technique enhances sensitivity to gravitational wave signals. From the definition of the single-arm response to the ALDM effect, the subtraction between the two channels cancels out these common-mode noises.
Working with the time derivative of the phase difference, rather than the phase difference itself, offers significant advantages for signal processing. It facilitates the implementation of TDI techniques and enable the separation of the ALDM signal from slowly varying systematic effects. Additionally, the frequency dependence of the signal provides a distinctive signature, aiding in the discrimination against potential background sources.
IV.2 Signals for different TDI combinations
@@
For each TDI combination, the ALDM-induced signal is weighted by a factor that depends on the arm length $L$ and the ALDM mass $m_{a}$. For example, the Sagnac variable $\alpha$ is given by Armstrong et al. (1999)
$\alpha(t)$ $=\Big{[}\bar{\eta}_{12}+D_{12}\bar{\eta}_{23}+D_{12}D_{23}\bar{\eta}_{31}\Big{]}-\Big{[}\eta_{13}+D_{13}\eta_{32}+D_{13}D_{32}\eta_{21}\Big{]}$ (27)
@@
where $D_{ij}$ represents the delay operator corresponding to the light travel time between S/C $i$ and $j$. Extending this to the second-generation Sagnac variable Vallisneri (2005), we obtain:
$\alpha_{2}(t)$ $=\Big{(}1-D_{12}D_{23}D_{31}\Big{)}\alpha$ (28)
@@
Unlike the modifications proposed in Gué et al. (2025); Yao et al. (2025), our approach not only leverages the Sagnac $\alpha$ combination but is also applicable to other TDI combinations. For example, the X variable reads
$X(t)$ $=\Big{(}1-D_{13}D_{31}\Big{)}\Big{(}\eta_{12}+D_{12}\eta_{21}\Big{)}-\Big{(}1-D_{12}D_{21}\Big{)}\Big{(}\bar{\eta}_{13}+D_{13}\bar{\eta}_{31}\Big{)}$ (29)
@@
The fully symmetric $\zeta$ variable is expressed as Muratore et al. (2022)
$\zeta(t)$ $=D_{12}\Big{(}\bar{\eta}_{31}-\eta_{32}\Big{)}+D_{23}\Big{(}\bar{\eta}_{12}-\eta_{13}\Big{)}+D_{31}\Big{(}\bar{\eta}_{23}-\eta_{21}\Big{)}$ (31)
@@
with a related second-generation variable Muratore et al. (2020)
$\zeta_{2}(t)$ $=\Big{(}D_{31}-D_{12}D_{23}\Big{)}\zeta$ (32)
$=-12\,\frac{g_{a\gamma}}{\omega_{r}}\sqrt{2\rho_{\mathrm{DM}}}\sin^{2}\frac{m_{a}L}{2}\cos\Big{(}m_{a}t+\Phi-3m_{a}L\Big{)}.$
@@
$S_{i}(f)=\frac{4C_{i}^{2}(f)g_{a\gamma}^{2}\rho_{\mathrm{DM}}}{\pi^{2}\omega_{r}^{2}T_{\mathrm{obs}}}\frac{\sin^{2}[\pi(f-f_{a})T_{\mathrm{obs}}]}{(f-f_{a})^{2}},$ (34)
@@
$C_{X}(f)=2\sin^{2}(2\pi fL),\quad C_{X_{2}}(f)=4\sin^{2}(2\pi fL)\sin(4\pi fL),$ (35)
$C_{\alpha}(f)=\sin(3\pi fL),\quad C_{\alpha_{2}}(f)=2\sin^{2}(3\pi fL),$ (36)
$C_{\zeta}(f)=3\sin(\pi fL),\quad C_{\zeta_{2}}(f)=6\sin^{2}(\pi fL).$ (37)
In the limit of $f\to f_{a}$, Eq. (34) simplifies to
$S_{i}(f)=4C_{i}^{2}(f)g_{a\gamma}^{2}\rho_{\mathrm{DM}}T_{\mathrm{obs}}/\omega_{r}^{2}.$ (38)
IV.3 Expected sensitivity
The two primary noise sources that limit the sensitivity of SGWDs are optical metrology system (OMS) noise and test mass acceleration noise. The OMS noise arises from imperfections in the laser interferometry system used to measure the distance between test masses in different spacecraft, such as residual frequency noise, shot noise, and pathlength noise. The acceleration noise originates from non-gravitational forces acting on the test masses. In our analysis, we assume that the proposed modifications to the OB do not significantly alter the noise PSDs. For completeness, we note that the noise PSDs for the $X$, $\alpha$, and $\zeta$ channels, as well as their second-generation counterparts, are given by expressions of the form Hartwig et al. (2023) [^2] Here we assume that the OMS and acceleration contributions for the Doppler variables $\eta$ and $\bar{\eta}$ are uncorrelated.
$N_{X}(f)$ $=16\sin^{2}(2\pi fL)\{[3+\cos(4\pi fL)]S_{\mathrm{acc}}(f)+S_{\mathrm{oms}}(f)\},$ (39)
$N_{X_{2}}(f)$ $=4\sin^{2}(4\pi fL)\,N_{X}(f),$ (40)
$N_{\alpha}(f)$ $=8\Big{[}2\sin^{2}(\pi fL)+\sin^{2}(3\pi fL)\Big{]}S_{\mathrm{acc}}(f)+6S_{\mathrm{oms}}(f),$ (41)
$N_{\alpha_{2}}(f)$ $=4\sin^{2}(3\pi fL)\,N_{\alpha}(f),$ (42)
$N_{\zeta}(f)$ $=6\Big{[}4\sin^{2}(\pi fL)S_{\mathrm{acc}}(f)+S_{\mathrm{oms}}(f)\Big{]},$ (43)
$N_{\zeta_{2}}(f)$ $=4\sin^{2}(\pi fL)\,N_{\zeta}(f).$ (44)
Here, $S_{\mathrm{oms}}(f)$ and $S_{\mathrm{acc}}(f)$ denote the PSDs of the OMS and acceleration noises, respectively. For LISA, TianQin and Taji, these noise PSDs are given by Babak et al. (2021)
$S_{\mathrm{oms}}(f)$ $=\left(2\pi fP_{\mathrm{oms}}/c\right)^{2}\left[1+\left(\frac{2\times 10^{-3}\>\mathrm{Hz}}{f}\right)^{4}\right]\>\mathrm{Hz}^{-1}\,,$ (45)

### v2

@@ 3 Proposed Modification
$\Delta\phi=\Delta\phi_{\mathrm{Tx}}+\Delta\phi_{\mathrm{Rx}}+\Delta\phi_{\mathrm{PS}}$ (3.12)
represents the phase shift relative to the ISI. Here, $\Delta\phi_{\mathrm{Tx}}$ and $\Delta\phi_{\mathrm{Rx}}$ are the phase differences arising from the path length differences of the transmitted light and received light, respectively. $\Delta\phi_{\mathrm{PS}}$ is the phase shift introduced by the phase shifter, which can be preset on Earth. The term $I_{1}$ is the power of the reference beam in the AUXI, and $p_{\mathrm{AUXI}}^{2}=p_{C}^{2}+p_{D}^{2}$ denotes the ratio of the received optical power to the reference beam when the incoming light is in a pure p-polarized state. The parameter $u_{\mathrm{AUXI}}^{2}=u_{C}^{2}+u_{D}^{2}$ denotes the ratio of the power of the backscattered light to the reference beam. Likewise, the power of the received light in the AUXI is modulated by the birefringence effect
$p_{\mathrm{AUXI}}^{\prime}=\sin\theta p_{\mathrm{AUXI}}.$ (3.13)
For simplicity, we assume $I_{1}=I_{0}$ and $p_{\mathrm{AUXI}}=p_{\mathrm{ISI}}$. In addition, the heterodyne signals are considered to be balanced [Carleton:68, Fleddermann:2017zna] with $p_{k}=p/\sqrt{2},\,k\in\{A,B,C,D\}$. Under these conditions, the combined AC components of the signal differences from both interferometers can be expressed as
$AC\{(S_{D}-S_{C})+(S_{B}-S_{A})\}$ (3.14)
$=$ $\sqrt{2}pI_{0}\Big[\cos\theta\sin(\Omega t+\phi)+\sin\theta\sin(\Omega t+\phi+\Delta\phi)$
[…]
$=$ $\sqrt{2}pI_{0}\Big[\sin(\Omega t+\phi+\theta)+\sin\theta[\sin(\Omega t+\phi+\Delta\phi)-\cos(\Omega t+\phi)]$
[…]
@@
$AC\{(S_{D}-S_{C})+(S_{B}-S_{A})\}$ (3.15)
$=$ $\sqrt{2}pI_{0}\Big[\sin(\Omega t+\phi+\theta)$
[…]
The choice of $\Delta\phi=\pi/2$ is a crucial, as it maximizes the sensitivity to the ALDM signal by ensuring that the auxiliary interferometer operates at the optimal phase difference relative to the main interferometer. This configuration converts the polarization angle into a phase shift relative to $\phi_{\mathrm{ISI}}$, which can be directly measured by standard phase detection techniques.
@@
$\phi_{\mathrm{AUXI}}\approx\phi+\theta+\pi/2+\phi_{b},$ (3.16)
@@
$\phi_{b}\approx\left[(u_{A}-u_{B})\cos\psi-(u_{C}-u_{D})\sin\tilde{\psi}\right]\left[1-(u_{A}-u_{B})\sin\psi-(u_{C}-u_{D})\cos\tilde{\psi}\right$]. (3.17)
The phase $\phi_{\mathrm{AUXI}}$ is the key measurand of the AUXI. By subtracting $\phi_{\mathrm{ISI}}$ from $\phi_{\mathrm{AUXI}}$, we can recover the polarization angle as expressed in Eq. (3.1). [^2] In a more realistic case, we can consider $I_{1}=f_{I}^{2}I_{0}$ and $p_{\mathrm{AUXI}}=f_{p}p_{\mathrm{ISI}}=f_{p}\frac{p}{\sqrt{2}}$, where $f_{I}$ and $f_{p}$ are determined by the reflectance and transmittance coefficients of the beam splitters. Then the phase difference would be $\phi_{\mathrm{AUXI}}-\phi_{\mathrm{ISI}}\simeq\arctan(\frac{1}{f_{I}f_{p}}\tan\theta)$. Then the change in the polarization angle induced by the ALDM for a single-link light can be obtained naturally.
In the current design of SGWDs, the misalignment polarization angle $\theta_{0}$ is calibrated to be 0. For the modified OB we proposed above, it would be extremely difficult to detect the p-polarized component, given the weak received laser power and the further suppression. To circumvent this issue, an intentional initial misalignment angle needs to be included from the outset, ensuring that both s- and p-components are simultaneously present and measurable. The observable quantity is the modulation of this preexisting polarization angle due to the ALDM-induced birefringence, rather than the appearance of an additional polarization component.
The total measurement noise in the SGWDs mainly comes from three sources: photon (shot) noise, residual laser frequency noise and spacecraft jitter noise. Photon noise dominates at higher frequencies ($1\,\mathrm{mHz}$); at lower frequencies ($\approx 0.1\,\mathrm{mHz}$), other noise sources dominate. The received optical power affects only the photon-noise-limited part of the spectrum. The noise spectral density due to photon noise is $S_{\mathrm{shot}}\propto\lambda/\sqrt{P_{\mathrm{rec}}}$, where $\lambda$ denotes the wavelength of the laser and $P_{\mathrm{rec}}$ denotes the received optical power. Within a 10–50% reduction of received optical power, SGWDs’ science performance would not be seriously affected. The high-frequency photon-noise-limited sensitivity would degrade slightly, but overall mission goals would remain achievable. Therefore, the introduction of the misalignment angle is acceptable and meaningful.
@@
4 Sensitivity to the ALDM-photon coupling
@@
4.1 Single-arm signal
By differentiating the phase difference induced by the ALDM-induced birefringence, we define the single-arm response as  [^3] The single-arm signal seems a little different from the equation in [Gue:2024txz], this originates from the typo of the definition of the reduced Planck energy in their paper, which is supposed to be $E_{P}=\sqrt{\hbar c/\kappa}\simeq 2.435\times 10^{18}\,\mathrm{GeV}$. With the correct definition, the factor $\sqrt{8\pi G}$ would be canceled out.
$\eta_{rs}$ $\equiv\frac{1}{\omega_{r}}\frac{d\left(\phi_{\mathrm{AUXI}}-\phi_{\mathrm{ISI}}\right)}{dt}\approx\frac{1}{\omega_{r}}\frac{d\Delta\phi_{a}}{dt}$ (4.1)
@@ 4.1 Single-arm signal
$\bar{\eta}_{rs}\equiv\frac{1}{\omega_{r}}\frac{d\left(\phi_{\mathrm{ISI}}-\phi_{\mathrm{AUXI}}\right)}{dt}\approx-\frac{1}{\omega_{r}}\frac{d\Delta\phi_{a}}{dt},$ (4.2)
where $\omega_{r}$ denotes the laser frequency, and the subscripts $r$ and $s$ represent the indices of the S/C that received and sent the light beam, respectively. The single-arm responses $\eta_{rs}$ and $\bar{\eta}_{rs}$ provide a direct measure of the time derivative of the ALDM-induced phase shift. In practice, SGWDs employ TDI to suppress noises, such as laser frequency noise and OB motion noise, by combining the signals from multiple interferometric arms with appropriate time delays [pub.1058615630, Tinto:1999yr, Dhurandhar:2001tct, Tinto:2020fcc]. This technique enhances sensitivity to gravitational wave signals. From the definition of the single-arm response to the ALDM effect, the subtraction between the two channels cancels out these common-mode noises.
Working with the time derivative of the phase difference, rather than the phase difference itself, offers significant advantages for signal processing. It facilitates the implementation of TDI techniques and enables the separation of the ALDM signal from slowly varying systematic effects. Additionally, the frequency dependence of the signal provides a distinctive signature, aiding in the discrimination against potential background sources.
4.2 Signals for different TDI combinations
@@
For each TDI combination, the ALDM-induced signal is weighted by a factor that depends on the arm length $L$ and the ALDM mass $m_{a}$. For example, the Sagnac variable $\alpha$ is given by [pub.1058615630]
$\alpha(t)$ $=\Big[\bar{\eta}_{12}+D_{12}\bar{\eta}_{23}+D_{12}D_{23}\bar{\eta}_{31}\Big]-\Big[\eta_{13}+D_{13}\eta_{32}+D_{13}D_{32}\eta_{21}\Big]$ (4.3)
@@
where $D_{ij}$ represents the delay operator corresponding to the light travel time between S/C $i$ and $j$. Extending this to the second-generation Sagnac variable [Vallisneri:2005ji], we obtain:
$\alpha_{2}(t)$ $=\Big(1-D_{12}D_{23}D_{31}\Big)\alpha$ (4.4)
@@
While the modifications in [Gue:2024txz, Yao:2024hap] can be applied to various TDI combinations (see, e.g., the Appendix of [Yao:2024hap] and Ref. [Liu:2025hwn] [^4] A related construction in the circular-polarization setup is also discussed in [Liu:2025hwn], where the ALDM response is formed from linear combinations rather than from two helicity-defined virtual paths.), the resulting sensitivity is maximized for Sagnac-type configurations due to the helicity–direction binding. In contrast, our proposed approach provides a distinct framework that avoids modifying the laser polarization state. This allows for optimized sensitivity performance across a broader range of TDI channels, maintaining robustness in sensitivity even in non-Sagnac configurations. For example, the X variable reads
$X(t)$ $=\Big(1-D_{13}D_{31}\Big)\Big(\eta_{12}+D_{12}\eta_{21}\Big)-\Big(1-D_{12}D_{21}\Big)\Big(\bar{\eta}_{13}+D_{13}\bar{\eta}_{31}\Big)$ (4.5)
@@
The fully symmetric $\zeta$ variable is expressed as [Muratore:2021uqj]
$\zeta(t)$ $=D_{12}\Big(\bar{\eta}_{31}-\eta_{32}\Big)+D_{23}\Big(\bar{\eta}_{12}-\eta_{13}\Big)+D_{31}\Big(\bar{\eta}_{23}-\eta_{21}\Big)$ (4.7)
@@
with a related second-generation variable [Muratore:2020mdf]
$\zeta_{2}(t)$ $=\Big(D_{31}-D_{12}D_{23}\Big)\zeta$ (4.8)
[…]
@@
$S_{i}(f)=\frac{4C_{i}^{2}(f)g_{a\gamma}^{2}\rho_{\mathrm{DM}}}{\pi^{2}\omega_{r}^{2}T_{\mathrm{obs}}}\frac{\sin^{2}[\pi(f-f_{a})T_{\mathrm{obs}}]}{(f-f_{a})^{2}},$ (4.10)
@@
$C_{X}(f)=2\sin^{2}(2\pi fL),\quad C_{X_{2}}(f)=4\sin^{2}(2\pi fL)\sin(4\pi fL),$ (4.11)
$C_{\alpha}(f)=\sin(3\pi fL),\quad C_{\alpha_{2}}(f)=2\sin^{2}(3\pi fL),$ (4.12)
$C_{\zeta}(f)=3\sin(\pi fL),\quad C_{\zeta_{2}}(f)=6\sin^{2}(\pi fL).$ (4.13)
In the limit of $f\to f_{a}$, Eq. (4.10) simplifies to
$S_{i}(f)=4C_{i}^{2}(f)g_{a\gamma}^{2}\rho_{\mathrm{DM}}T_{\mathrm{obs}}/\omega_{r}^{2}.$ (4.14)
4.3 Expected sensitivity
The two primary noise sources that limit the sensitivity of SGWDs are optical metrology system (OMS) noise and test mass acceleration noise. The OMS noise arises from imperfections in the laser interferometry system used to measure the distance between test masses in different spacecraft, such as residual frequency noise, shot noise, and pathlength noise. The acceleration noise originates from non-gravitational forces acting on the test masses. Both are broadband and stochastic in nature.
Scattered-light noise, originating from parasitic reflections and microscopic surface scattering, is largely incoherent over time: while it interferes coherently within a single optical path, its phase rapidly decorrelates on time scales shorter than a few seconds due to thermal and mechanical perturbations. As a result, its contribution to the long-term phase noise power spectral density remains far below that of the primary noise sources. Provided that the optical layout is not modified in a way that enhances coherent low-frequency scattering paths, such scattered-light noise can be safely neglected in sensitivity estimates dominated by acceleration and readout noise.
In contrast, the ALDM-induced birefringence signal is a temporally coherent, narrowband effect, determined by the difference of the axion field amplitude between the emission and reception points of the laser beam. For the typical SGWDs arm length, the induced polarization rotation is of order $10^{-7}\,\mathrm{rad}$. Although the instantaneous amplitude of this signal is below the phase noise level, its long coherence time enables coherent integration that enhances the signal-to-noise ratio as $T^{1/2}$. Under these conditions, broadband random noise—including residual scattered-light fluctuations—averages down efficiently, while only narrowband or persistent scattering features near the axion Compton frequency could produce correlated artifacts.
In our analysis, we assume that the proposed modifications to the OB do not significantly alter the noise PSDs. For completeness, we note that the noise PSDs for the $X$, $\alpha$, and $\zeta$ channels, as well as their second-generation counterparts, are given by expressions of the form [Hartwig:2023pft] [^5] Here we assume that the OMS and acceleration contributions for the Doppler variables $\eta$ and $\bar{\eta}$ are uncorrelated.
$N_{X}(f)$ $=16\sin^{2}(2\pi fL)\{[3+\cos(4\pi fL)]S_{\mathrm{acc}}(f)+S_{\mathrm{oms}}(f)\},$ (4.15)
$N_{X_{2}}(f)$ $=4\sin^{2}(4\pi fL)\,N_{X}(f),$ (4.16)
$N_{\alpha}(f)$ $=8\Big[2\sin^{2}(\pi fL)+\sin^{2}(3\pi fL)\Big]S_{\mathrm{acc}}(f)+6S_{\mathrm{oms}}(f),$ (4.17)
$N_{\alpha_{2}}(f)$ $=4\sin^{2}(3\pi fL)\,N_{\alpha}(f),$ (4.18)
$N_{\zeta}(f)$ $=6\Big[4\sin^{2}(\pi fL)S_{\mathrm{acc}}(f)+S_{\mathrm{oms}}(f)\Big],$ (4.19)
$N_{\zeta_{2}}(f)$ $=4\sin^{2}(\pi fL)\,N_{\zeta}(f).$ (4.20)
Here, $S_{\mathrm{oms}}(f)$ and $S_{\mathrm{acc}}(f)$ denote the PSDs of the OMS and acceleration noises, respectively. For LISA, TianQin and Taji, these noise PSDs are given by [Babak:2021mhe]
$S_{\mathrm{oms}}(f)$ $=\left(2\pi fP_{\mathrm{oms}}/c\right)^{2}\left[1+\left(\frac{2\times 10^{-3}\>\mathrm{Hz}}{f}\right)^{4}\right]\>\mathrm{Hz}^{-1}\,$, (4.21)